OpenStudy (anonymous):

I NEED HELP SO BAD I WILL MEDAL FAN AND WRITE TESTIMONIAL!!!!! QUESTION IN COMMENTS

OpenStudy (anonymous):

A ball is thrown into the air with an upward velocity of 32 feet per second. Its height, h, in feet after t seconds is given by the function h(t) = –16t² + 32t + 6. What is the ball’s maximum height? How long does it take the ball to reach its maximum height? Round to the nearest hundredth, if necessary. (1 point) Reaches a maximum height of 22 feet after 1.00 second. Reaches a maximum height of 22 feet after 2.00 seconds. Reaches a maximum height of 44 feet after 2.17 seconds. Reaches a maximum height of 11 feet after 2.17 seconds.

OpenStudy (anonymous):

The maximum height of the ball, is when velocity is 0 have you done calculus? differentiation or no?

OpenStudy (anonymous):

This is algebra 1b so no calc

OpenStudy (anonymous):

ok. well not sure how to get you to a function for velocity. Other than to just tell you that velocity = -16t + 32 at the heighest point velocity is 0 (as the ball turns around to come back home) so 0 = -16t + 32 solve that for t t = 32/16 t = 2

OpenStudy (anonymous):

then whack that value of time in for t in your other equation

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

but you can kind of see why velocity would be that. velocity = distance/ time = h(t)/t and we just ignore the 6/t you get at the end, because the 6 is the initial height and that is irrespective of velocity

OpenStudy (anonymous):

-2?